Tin exists in Earth’s crust as SnO2. Calculate the percent composition by mass of Sn and O in SnO2.
1
Expert's answer
2020-10-30T04:03:21-0400

Molar mass of SnO2 = 213.71 g/mol


Mass of Sn present in one SnO2 molicule

= Atomic mass of Sn

= 181.71 g/mol


So,

percentage of Sn by mass in SnO2

= (Mass of Sn in SnO2/ Molar mass of SnO2)×100%

= (181.71/213.71) × 100 %

= 85.026 %



Atomic mass of oxygen (O) = 16 g/mol


total mass of oxygen present in SnO2

= [Atomic mass of oxygen (O) × 2] g/mol

= (16×2) g/mol

= 32 g/mol


So,

Percentage of Oxygen by mass in SnO2

= (Mass of oxygen in SnO2/total mass of SnO2)×100 %

= (32/213.71)×100 %

= 14.973 %



Hence,

SnO2 composed by 14.973 % of oxygen and 85.026 % of Sn by mass